# Energy-time and time-bin entanglement: past, present and future

### v1

@@ I Introduction
Franson proposed a novel setup for generating and analyzing a radically different form of entanglement in 1989 [16]. It exploited the random nature in the generation time of a photon pair from the cascaded decay in a three-level atomic system, which was then analyzed by unbalanced Mach-Zehnder interferometers. The scheme was demonstrated already shortly afterwards [17, 18] and it quickly became the mainstay for long-distance Bell violation experiments during the 90s [19, 20].
@@ II Energy-time entanglement and early experiments
In 1991, Brendel et al. demonstrated the first two-photon high-visibility experiment using Franson interferometers, by employing a short coincidence window in the detection electronics post-selecting only the overlapping indistinguhishable $|S_{a}S_{b}\rangle$ and $|L_{a}L_{b}\rangle$ components (Figs. 2g and 2h) [43]. This post-selection procedure does recover a high-visibility Bell inequality violation usable as an entanglement test, but has the downside that it removes the direct link between Bell inequality violations and violations of local realism [44]; we will return to how to restore this link in Section III. Nonetheless, the benefits of the post-selection procedure caused it to quickly become adopted, which led to the first energy-time Bell inequality violations (Figs. 2i and 2j) [45, 46] and also the first use of an optical fiber interferometer two-photon energy-time experiments (Figs. 2k and 2l) [47], paving the way for what would come later in terms of long-distance Bell inequality violations.
@@ V Integrated sources of energy-time and time-bin entanglement
The use of a waveguide structure within a periodically poled bulk crystal for the efficient generation of energy-time and time-bin entanglement has since then been used in many demonstrations [110, 111, 112, 113, 114, 115, 116, 117, 55, 69, 72]. More recently new advances in thin-film lithium niobate technology, where the waveguide is considerably shallower, allowed considerable gains in generation efficiency, as well as ultra-broad band emission [118, 119, 120, 121] (Fig. 7a).
@@
Further work was done in optimizing time-bin photon pair extraction from the quantum dot [167] as well as from a quantum dot embedded in a nanowire [168]. In [169], Franson interferometry was used to characterize the inhomogeneous broadening of the bi-exciton state in a quantum dot, although the visibility was far too low to violate a Bell inequality. Later, a successful energy-time Bell violation was shown directly from a quantum dot emission [170] (Fig. 7e), the only result of this type so far from this type of source.
@@ VI Applications
Another example is superdense coding that was performed over short optical fiber links, together with complete Bell state measurements relying on hyperentanglement between polarization and energy-time [214]. Finally, entanglement purification of has been successfully performed, once again relying on polarization/energy-time hyperentanglement, but in this case to realize the required CNOT operation as well as lowering the requirement to only one copy, thus requiring only one photon pair [32] (Fig. 8d).
@@
A major difficulty expanding to higher-dimensional time-bin and energy-time entanglement is the difficulty in building stable multi-arm interferometers. One solution is to employ cascaded two-arm interferometers with varying delays [74, 226]. One study showed that such cascaded interferometers are important for the security of high-dimensional QKD protocols based on energy-time entanglement [227]. Another study showed that high-dimensional entanglement can be characterized using only two output interferometers, applying the technique for time-bin entanglement [228].
@@
This concept applied to energy-time entanglement was used to demonstrate one of the first high-dimensional QKD experiments [229]. A security proof was then derived employing an analogy to continuous-variable (CV) entanglement, when dual-basis interferometry is used, alternating between Franson’s interferometers and its frequency conjugate [230]. A more practical proof considering decoy states to tackle multi-pair emission is presented in [231], where a variation of a setup generating energy-time entanglement is employed where the measurements are performed in the conjugate time and frequency bases.
An experimental implementation for high-dimensional energy-time entanglement QKD is then carried out where the time basis is used to generate the correlated key bits with a dimensionality of 1000, and the energy basis with Franson interferometers is used to check security [232]. A very recent demonstration combined an SFWM source in integrated photonics, with the analyzers also implemented in photonic chips with unbalanced cascaded Mach-Zehnder interferometers generating up to 8-dimensional time-bin entanglement, and entanglement-based QKD over up to 60 km propagation over optical fibers [233].
@@ VIII Acknowledgements
This work is part of the project SECuRe quantum communication based on Energy-Time entanglement (SECRET), QuantERA Call 2019. We wish to thank Zenith Linköping University, the QuantERA grant SECRET (VR 2019-00392), VR regular grant (VR 2023-05031), the EU-funded project FoQaCiA, the MCINN/AEI (Project No. PID2020-113738GB-I00) for financial support.

### v2

@@ I Introduction
Franson proposed a novel setup for generating and analyzing a radically different form of entanglement in 1989 [16]. It exploited the random nature in the generation time of a photon pair from the cascaded decay in a three-level atomic system, which was then analyzed by unbalanced Mach-Zehnder interferometers [17]. The scheme was demonstrated already shortly afterwards [18, 19] and it quickly became the mainstay for long-distance Bell violation experiments during the 90s [20, 21].
@@ II Energy-time entanglement and early experiments
In 1991, Brendel et al. demonstrated the first two-photon high-visibility experiment using Franson interferometers, by employing a short coincidence window in the detection electronics post-selecting only the overlapping indistinguhishable $|S_{a}S_{b}\rangle$ and $|L_{a}L_{b}\rangle$ components (Figs. 2g and 2h) [44]. This post-selection procedure does recover a high-visibility Bell inequality violation usable as an entanglement test, but has the downside that it removes the direct link between Bell inequality violations and violations of local realism [45]; we will return to how to restore this link in Section III. Nonetheless, the benefits of the post-selection procedure caused it to quickly become adopted, which led to the first energy-time Bell inequality violations (Figs. 2i and 2j) [46, 47] and also the first use of an optical fiber interferometer two-photon energy-time experiments (Figs. 2k and 2l) [48], paving the way for what would come later in terms of long-distance Bell inequality violations. It is also worth to notice that unbalanced interferometers have also been extensively employed to measure time-bin states in prepare-and-measure scenarios, such as in [49, 50]
@@ V Integrated sources of energy-time and time-bin entanglement
The use of a waveguide structure within a periodically poled bulk crystal for the efficient generation of energy-time and time-bin entanglement has since then been used in many demonstrations [113, 114, 115, 116, 117, 118, 119, 120, 58, 72, 75, 121, 122]. More recently new advances in thin-film lithium niobate technology, where the waveguide is considerably shallower, allowed considerable gains in generation efficiency, as well as ultra-broad band emission [123, 124, 125, 126] (Fig. 7a). One demonstrated technique that is relevant for ultra-broad band emission is the measurement of geometric phases due to their non-dispersive nature [127].
@@
Further work was done in optimizing time-bin photon pair extraction from the quantum dot [173] as well as from a quantum dot embedded in a nanowire [174]. In [175], Franson interferometry was used to characterize the inhomogeneous broadening of the bi-exciton state in a quantum dot, although the visibility was far too low to violate a Bell inequality. Later, a successful energy-time Bell violation was shown directly from a quantum dot emission [176] (Fig. 7e), the only result of this type so far from this type of source. Finally, energy-time entanglement was also generated from inelastic scattering off a quantum dot coupled to a waveguide [177] and from a Rubidium atom coupled to a fiber Fabry-Pérot cavity [178].
@@ VI Applications
Another example is superdense coding that was performed over short optical fiber links, together with complete Bell state measurements relying on hyperentanglement between polarization and energy-time [222]. Finally, entanglement purification [223, 224, 225] has been successfully performed, once again relying on polarization/energy-time hyperentanglement, but in this case to realize the required CNOT operation as well as lowering the requirement to only one copy, thus requiring only one photon pair [33] (Fig. 8d).
@@
A major difficulty expanding to higher-dimensional time-bin and energy-time entanglement is the difficulty in building stable multi-arm interferometers. One solution is to employ cascaded two-arm interferometers with varying delays [77, 237]. One study showed that such cascaded interferometers are important for the security of high-dimensional QKD protocols based on energy-time entanglement [238]. Another study showed that high-dimensional entanglement can be characterized using only two output interferometers, applying the technique for time-bin entanglement [239]. Recently another approach is based on the use of multi-mode fibres to carry out generalised projective measurements for high-dimensional time-bin entanglement [240].
@@
This concept applied to energy-time entanglement was used to demonstrate one of the first high-dimensional QKD experiments [241], with a more recent experiment also demonstrating quantum steering [242]. A security proof was then derived employing an analogy to continuous-variable (CV) entanglement, when dual-basis interferometry is used, alternating between Franson’s interferometers and its frequency conjugate [243]. A more practical proof considering decoy states to tackle multi-pair emission is presented in [244], where a variation of a setup generating energy-time entanglement is employed where the measurements are performed in the conjugate time and frequency bases.
An experimental implementation for high-dimensional energy-time entanglement QKD is then carried out where the time basis is used to generate the correlated key bits with a dimensionality of 1000, and the energy basis with Franson interferometers is used to check security [245]. A very recent demonstration combined an SFWM source in integrated photonics, with the analyzers also implemented in photonic chips with unbalanced cascaded Mach-Zehnder interferometers generating up to 8-dimensional time-bin entanglement, and entanglement-based QKD over up to 60 km propagation over optical fibers [246]. Finally, recent efforts have also focused on the characterization of spectral Franson interference for high-dimensional applications [247].
@@
In spite of its difficulties, tri-partite energy-time entanglement was actually generated by taking advantage of two SPDC cascaded processes, where one of the generated photons in the first process acts as a pump in the second [256]. This technique was then successfully used to demonstrate three-photon non-local interference with the individual propagation paths all placed in the same unbalanced Michelson interferometer for extra stability [257]. In spite of the low creation of three-photon events, this technique has the advantage of not needing to post-select a three-photon state out of a four photon state produced by coherent SPDC processes.
@@ VIII Acknowledgements
The authors would like to thank Yu-Bo Sheng, Gui-Lu Long, Jian Wang, Michael Lubasch, Kai-Chi Chang, Mehul Malik and Zi-Qi Zeng for helpful discussions. This work is part of the project SECuRe quantum communication based on Energy-Time entanglement (SECRET), QuantERA Call 2019. We wish to thank Zenith Linköping University, the QuantERA grant SECRET (VR 2019-00392), VR regular grant (VR 2023-05031), the EU-funded project FoQaCiA, the MCINN/AEI (Project No. PID2020-113738GB-I00) for financial support.